# Photonic Exceptional Points in Holography and QCD

### v4

@@ 4.2 Inhomogeneous case: holographic lattice for three-site photonic EP
For deeper lattices with three sites, one could extend this model to inhomogeneous $s(x^{1})$, and with RG flows to a PT-symmetric IR fixed point, where non-Hermitian “cooling” which is the decrease in entropy, occurs via the horizon shrinkage.
@@
Then, in the PT-unbroken phase, where the gain/loss satisfies the inequality $\Gamma<\Gamma_{c}(k)$, and by photonic analogy $\Gamma_{c}\sim g\sqrt{3}$, the QNMs $\omega_{n}(k)$ are real for each band, and the dispersion is $\omega(k)\approx\omega_{0}+2g\cos(k\ell)$. The bulk profile $\phi(z,x)$ is a smooth modulation with no instabilities. The boundary Green’s function $G(x,x^{\prime};\omega)=\langle\mathcal{O}(x)\mathcal{O}^{\dagger}(x^{\prime})\rangle$ decays exponentially for $|x-x^{\prime}|>\ell$, which mimics the localized resonator modes.
@@
Then, in the PT-broken phase, $\Gamma>\Gamma_{c}$, the complex conjugate branches are $\omega_{\pm}(k)=\omega_{r}(k)\pm i\omega_{i}(k)$, with $\omega_{i}>0$ for amplification, where we have gain-dominated bands. In the bulk, we get instability via the complex field $\phi(z,x)$, which has growing modes along $x$ with Petrov-type shifts. Considering the full backreaction, the metric warps into a striped geometry where $g_{xx}(z,x)\sim 1+\delta\cos(2\pi x/a)$, and the NEC is broken locally, akin to holographic superconductors with modulated order.
@@ 4.4 Holographic phase rigidity and Petermann factors around EPs
which quantifies the complexness of the wave functions in PT-broken symmetry systems. Note that around an EP, the Petermann factor diverges and the phase rigidities vanish. Also, around the EPs, as the vector space becomes severely skewed, we have a losing dimensionality effect, which is related to the divergence of the Petermann factor.
@@
Next, in figure 14, the Petermann factor and phase rigidities as a function of gain/loss $\gamma$ are shown while other parameters are fixed. Here the “dip-spike” structure around $\gamma\approx 0.8-1.0$ is where the EP occurs. The minsep which are the dashed black curve, reaches a minimum indicates coalescing eigenvalues and phase rigidity $r_{n}$ varies strongly near the EP.
@@ 5 EPs, photonic molecules and quark bound states
An interesting model that one can use is the picture of the “bound states” for a ternary, parity-time-symmetric photonic laser molecule. For instance, mesons, which consist of a quark-antiquark pair, can model the EP2 structures, and baryons, which consist of three color quarks, can be associated with EP3.
As noted in [42], by using the ternary systems and then breaking the PT symmetry and “perturbing” the anti-PT symmetry, a pure single-mode spectrum can be achieved. This is similar to the confinement/deconfinement phase transition. So we can relate these two matrices as
@@
where $\Sigma_{0}$ maximizes the static Lagrangian and $T^{a}$ are the $SU(3)$ generators. Here, $\Sigma$ describes the pseudoscalar meson octet
@@ 6 Exceptional point and timelike entanglement entropy
As noted EPs mark a transition where the non-Hermitian Hamiltonian’s spectrum shifts from real PT-unbroken and stable modes to complex conjugate pairs with PT-broken and amplification or decay. This manifests in temporal dynamics as exponential growth or decay, which TEE is uniquely suited to capture. Note that in the PT-broken phase, our photonic system’s modes have $\text{Im}\neq 0$, leading to non-unitary time evolution. The imaginary part of TEE, which can probe these temporal commutators, is given by
@@
where $\beta=t_{R}-t_{L}$ is the boundary length, $s=\gamma+i\ \Gamma$ is the source, $\epsilon$ is a UV cutoff, and $\Theta$ is the step function signaling PT breaking. So when $\Gamma<\Gamma_{c}$, $S_{T}^{I}\approx 0$, as there would be no dissipation in this case. At $\Gamma=\Gamma_{c}$, we get an EP, as there is coalescence where multiple bulk surfaces compete and $S_{T}$ develops a $\sqrt{\Gamma-\Gamma_{c}}$ branch point. When $\Gamma>\Gamma_{c}$, we get a PT-broken phase where $S_{T}^{I}\neq 0\sim|\text{Im}(\omega)|\beta$, where for positive imaginary part there is amplification and for negative part there is decay. There is a discontinuity in the imaginary part which signals the transition.
@@
As for the experimental setup, for a ring cavity that operates at an exceptional point with PT-symmetric refractive index modulation [81],
@@ 8 Conclusion
In this note, based on the behavior of fidelity susceptibility, chiral symmetry breaking, and the number of spectra and coalescing levels, we proposed the holographic dual of optical exceptional points to be the end-wall in holographic confining geometries, and based on the order of EPs and their numbers in the system, we connected them with the hard-wall or soft-wall models. We also discussed the effects of both the EPs and end-wall on the chaos. Specifically, for different patterns of gain and loss we discussed the holographic spectra and showed that they match the behaviors coming from optical experiments. For symmetric pumping pattern in the ternary coupled microcavity model like the anti-PT sweep, we built the eigenvalue trajectories, radial profiles of the dominant lasing eigenmode, and also showed the behavior of growth rates of modes. We also built the phase diagram distinguishing unbroken PT versus broken PT phase.

### v5

@@ 4.2 Inhomogeneous case: holographic lattice for three-site photonic EP
For deeper lattices with three sites, one could extend this model to inhomogeneous $s(x^{1})$, and with RG flows to a $\mathcal{PT}$-symmetric IR fixed point, where non-Hermitian “cooling” which is the decrease in entropy, occurs via the horizon shrinkage.
@@
Then, in the $\mathcal{PT}$-unbroken phase, where the gain/loss satisfies the inequality $\Gamma<\Gamma_{c}(k)$, and by photonic analogy $\Gamma_{c}\sim g\sqrt{3}$, the QNMs $\omega_{n}(k)$ are real for each band, and the dispersion is $\omega(k)\approx\omega_{0}+2g\cos(k\ell)$. The bulk profile $\phi(z,x)$ is a smooth modulation with no instabilities. The boundary Green’s function $G(x,x^{\prime};\omega)=\langle\mathcal{O}(x)\mathcal{O}^{\dagger}(x^{\prime})\rangle$ decays exponentially for $|x-x^{\prime}|>\ell$, which mimics the localized resonator modes.
@@
Then, in the $\mathcal{PT}$-broken phase, $\Gamma>\Gamma_{c}$, the complex conjugate branches are $\omega_{\pm}(k)=\omega_{r}(k)\pm i\omega_{i}(k)$, with $\omega_{i}>0$ for amplification, where we have gain-dominated bands. In the bulk, we get instability via the complex field $\phi(z,x)$, which has growing modes along $x$ with Petrov-type shifts. Considering the full backreaction, the metric warps into a striped geometry where $g_{xx}(z,x)\sim 1+\delta\cos(2\pi x/a)$, and the NEC is broken locally, akin to holographic superconductors with modulated order.
@@ 4.4 Holographic phase rigidity and Petermann factors around EPs
which quantifies the complexness of the wave functions in $\mathcal{PT}$-broken symmetry systems. Note that around an EP, the Petermann factor diverges and the phase rigidities vanish. Also, around the EPs, as the vector space becomes severely skewed, we have a losing dimensionality effect, which is related to the divergence of the Petermann factor.
@@
Next, in figure 14, the Petermann factor and phase rigidities as a function of gain/loss $\gamma$ are shown while other parameters are fixed. Here the “dip-spike” structure around $\gamma\approx 0.8-1.0$ is where the EP occurs. The minseps which are the dashed black curve, reaches a minimum indicates coalescing eigenvalues and phase rigidity $r_{n}$ varies strongly near the EP.
@@ 5 EPs, photonic molecules and quark bound states
An interesting model that one can use is the picture of the bound states for a ternary, parity-time-symmetric photonic laser molecule. For instance, mesons, which consist of a quark-antiquark pair, can model the EP2 structures, and baryons, which consist of three color quarks, can be associated with EP3. However, note that this section is only an “analogy” and unlike previous section, we don’t make any holography argument here and only are building a toy model.
So as noted in [25], by using the ternary systems and then breaking the PT symmetry and perturbing the anti-PT symmetry, a pure single-mode spectrum can be achieved. This is similar to the confinement/deconfinement phase transition. So by an analogy, we can relate these two matrices as
@@
where $\Sigma_{0}$ maximizes the static Lagrangian and $T^{a}$ are the $SU(3)$ generators. Here, $\Sigma$ describes the pseudo-scalar meson octet
@@ 6 Exceptional point and timelike entanglement entropy
As noted EPs mark a transition where the non-Hermitian Hamiltonian’s spectrum shifts from real $\mathcal{PT}$-unbroken and stable modes to complex conjugate pairs with $\mathcal{PT}$-broken and amplification or decay. This manifests in temporal dynamics as exponential growth or decay, which TEE is uniquely suited to capture. Note that in the $\mathcal{PT}$-broken phase, our photonic system’s modes have $\text{Im}\neq 0$, leading to non-unitary time evolution. The imaginary part of TEE, which can probe these temporal commutators, is given by
@@
where $\beta=t_{R}-t_{L}$ is the boundary length, $s=\gamma+i\ \Gamma$ is the source, $\epsilon$ is a UV cutoff, and $\Theta$ is the step function signaling $\mathcal{PT}$ breaking. So when $\Gamma<\Gamma_{c}$, $S_{T}^{I}\approx 0$, as there would be no dissipation in this case. At $\Gamma=\Gamma_{c}$, we get an EP, as there is coalescence where multiple bulk surfaces compete and $S_{T}$ develops a $\sqrt{\Gamma-\Gamma_{c}}$ branch point. When $\Gamma>\Gamma_{c}$, we get a $\mathcal{PT}$-broken phase where $S_{T}^{I}\neq 0\sim|\text{Im}(\omega)|\beta$, where for positive imaginary part there is amplification and for negative part there is decay. There is a discontinuity in the imaginary part which signals the transition.
@@
As for the experimental setup, for a ring cavity that operates at an exceptional point with $\mathcal{PT}$-symmetric refractive index modulation [62],
@@
Another interesting point worths to mention here is the recent consideration of an “observer” in a form of a clock [67, 68], a charged black hole in de Sitter space [69], or two coupled SYK systems [70, 71], which can mimic the properties of an open quantum system and to explain properties such as complex phase space in the one-loop gravitational path integral. These observers, by continuous monitoring can enhance chaos in the systems similar to EPs and can cancel the complex phases.
@@ 7.1 Numerical investigation of EPs in $\theta$-vacuum QCD model
7.2 Controlled non-Hermitian deformation of $\theta$-vacuum toy model
In the $\theta$-vacuum model, we know that at $\theta\approx\pi$, the effective theory has two nearly degenerate topological branches.
From the action 7.14, the second term, which is the $\theta$-term becomes a topological term as
$\frac{\theta}{4\pi}\int_{0}^{\beta}d\tau\ (\varphi^{*}\dot{\varphi}-\dot{\varphi}^{*}\varphi)=i\theta\ \nu,$ (7.28)
where $\nu$ is the winding number and $\nu\in\mathbb{Z}$.
Also, from the action, the extended Hamiltonian would be
$H=\frac{1}{2m}\Pi_{r}^{2}+\frac{1}{2mr^{2}}\left(\Pi_{\vartheta}-\frac{\theta+i\epsilon}{2\pi}r^{2}\right)^{2}+\frac{g}{4}(r^{2}-1)^{2},$ (7.29)
where the canonical momenta are
$\Pi_{r}=m\dot{r},\ \ \ \ \ \ \ \ \Pi_{\vartheta}=mr^{2}\dot{\vartheta}+\frac{\theta}{2\pi}r^{2}.$ (7.30)
Then, we can make $\theta$ complex as $\theta\to\theta+i\epsilon$, and then similar to works such as [70, 71], we can study Liouvillian, dissipative gap and the spectrum of the deformed operator as a function of $\epsilon$.
At low energy, $r$ is localized near $r=1$, so $\frac{1}{r^{2}}\to 1$, and the radial excitations decouple, and the angular energy levels would be
$E_{n}(\epsilon)=\frac{1}{2m}\left(n-\frac{\theta+i\epsilon}{2\pi}\right).$ (7.31)
Near $\theta=\pi$, we have degeneracy, and can expand around it by defining $\delta\theta=\theta-\pi$.
Then, the effective Hamiltonian would be
$H_{\text{eff}}=E_{0}\mathbb{I}+\begin{pmatrix}\alpha\ \delta\theta+i\gamma\epsilon&i\kappa\epsilon\\ i\kappa\epsilon&-\alpha\ \delta\theta-i\gamma\epsilon\end{pmatrix}$, (7.32)
where $\alpha\sim\frac{1}{2\pi m}$, and $\gamma$, $\kappa$ are real constants fixed by the angular sector.
In fact, if we consider the exact deformed Hamiltonian, we could write
[…]
which
$\gamma_{n}=\frac{1}{2m\pi r^{2}}\left(n-\frac{\theta}{2\pi}\right).$ (7.34)
So $\gamma$ is the rate at which the energy of each angular sector moves into the complex plane. Also, $\kappa$ is a second order induced non-Hermitian mixing, coming from the projection onto finite subspace, integrating out higher angular modes or radial excitations, or allowing weak environmental coupling. So we can write
[…]
This parameter actually comes from virtual transitions through eliminated states and is zero if the system is closed. Note that in low energy and in subspace $r\to 1$.
The eigenvalues of the Hamiltonian are
$\lambda_{\pm}=E_{0}\pm\sqrt{\alpha^{2}\delta\theta^{2}-(\kappa^{2}-\gamma^{2})\epsilon^{2}+2i\ \alpha\gamma\delta\theta\ \epsilon},$ (7.36)
which for the EP, gives the condition of
$\alpha^{2}\delta\theta^{2}=(\kappa^{2}-\gamma^{2})\epsilon^{2},$ (7.37)
where $\delta\theta$, and $\epsilon$ are real. At these points, the eigenvalues and eigenvectors coalesce, the Hamiltonian becomes non-diagonalizable and the spectrum exhibits a square-root branch point.
The Liouvillian, $\frac{d\rho}{dt}=\mathcal{L}(\rho)$, using the Lindblad formalism would be
$\mathcal{L}=-i[H_{\theta},\rho]+\sum_{\mu}\Big(L_{\mu}\rho L^{\dagger}_{\mu}-\frac{1}{2}\{L^{\dagger}_{\mu}L_{\mu},\rho\}\Big).$ (7.38)
For our the $\theta$-term model, as the angular degree of freedom is the main structure, a minimal, controlled choice for $L$ would be
$L=\sqrt{\varepsilon}\ \left(\Pi_{\vartheta}-\frac{\theta}{2\pi}\right).$ (7.39)
The dissipative gap then would be
$\Delta_{\text{diss}}=\frac{\varepsilon}{2}\left(1-\frac{\theta}{\pi}\right)^{2},$ (7.40)
which closes near $\theta=\pi$.
If we define $A\equiv\Pi_{\vartheta}-\theta/2\pi$, the Liouvillian could be written as
$\mathcal{L}\rho=-i[H_{\theta},\rho]+\varepsilon\left[A\rho A-\frac{1}{2}\{A^{2},\rho\}\right],$ (7.41)
and in the angular momentum eigenbasis we have
[…]
Then, for $n\neq m$, we have
$\frac{d\rho_{nm}}{dt}=\left[-i(E_{n}-E_{m})-\frac{\varepsilon}{2}(a_{n}-a_{m})^{2}\right]\rho_{nm},$ (7.43)
and therefore, the Liouvillian eigenvalues are
$\lambda_{nm}=-\frac{\varepsilon}{2}(a_{n}-a_{m})^{2}-i(E_{n}-E_{m}).$ (7.44)
The retarded Greens function then would be
$G^{R}(\omega)=\frac{\omega+(\Delta-i\varepsilon\delta\gamma)\sigma_{z}+i\varepsilon\kappa\sigma_{x}}{\omega^{2}-(\Delta-i\epsilon\delta\gamma)^{2}+(\varepsilon\kappa)^{2}},$ (7.45)
where its poles are at
$\omega_{\pm}=\pm\sqrt{(\Delta-i\varepsilon\delta\gamma)^{2}-(\varepsilon\kappa)^{2}}.$ (7.46)
Based on the Liouvillian eigenvalues, one can decompose the Green’s function as
$\mathcal{G}(\omega)=\sum_{n\neq m}\frac{|nm)(nm|}{\omega+\frac{\varepsilon}{2}(a_{n}-a_{m})^{2}+i(E_{n}-E_{m})},$ (7.47)
where $||nm)\equiv\ket{n}\bra{m}$ are coherence modes which diagonalizes the Liouvillian superoperator, and $a_{n}\equiv n-\frac{\theta}{2\pi}$ is the $\theta$-shifted angular momentum eigenvalue.
From this Green’s function, the closest pole to the origin would be
$\Delta_{\text{diss}}=\min\limits_{n\neq m}\frac{\varepsilon}{2}(a_{n}-a_{m})^{2}=\frac{\varepsilon}{2}\left(1-\frac{\theta}{\pi}\right)^{2},$ (7.48)
where it becomes long-lived at $\theta=\pi$, and signals dissipative criticality there.
@@ 8 Conclusion
In this note, based on the behavior of fidelity susceptibility, chiral symmetry breaking, and the number of spectra and coalescing levels, we proposed the holographic dual of optical exceptional points to be the end-wall in holographic confining geometries, and based on the order of EPs and their numbers in the system, we connected them with the hard-wall or soft-wall models. We also discussed the effects of both the EPs and end-wall on the chaos. Specifically, for different patterns of gain and loss we discussed the holographic spectra and showed that they match the behaviors coming from optical experiments. For symmetric pumping pattern in the ternary coupled microcavity model like the anti-$\mathcal{PT}$ sweep, we built the eigenvalue trajectories, radial profiles of the dominant lasing eigenmode, and also showed the behavior of growth rates of modes. We also built the phase diagram distinguishing unbroken $\mathcal{PT}$ versus broken $\mathcal{PT}$ phase.